OpenStudy (olsolneo123):

Which of the following equations represents a line that is perpendicular to the line that passes through the points below?

OpenStudy (solomonzelman):

perpendicular line slope: Comparing to the (m) the slope of of a regular line, perpendicular is going to be -1/m OR multiplicative inverse times -1. Regular slope, \[slope=\frac{y_1-y_2}{x_1-x_2}\]

OpenStudy (olsolneo123):

\[\left( -2,-8\right)\] \[\left( 0,-2 \right)\] those are my points. here are the answer choices \[y=-3x+5\] \[y=3x+5\] \[y=\left(\begin{matrix}1 \\ 3\end{matrix}\right)x+5\] \[y=-\left(\begin{matrix}1 \\ 3\end{matrix}\right)x+5\]

OpenStudy (solomonzelman):

\[\huge\color{blue}{\frac{y_1-y_2}{x_1-x_2}}\]\[\huge\color{blue}{\frac{(-8)-(-2)}{(-2)-(0)}}~~~->~~~\huge\color{blue}{\frac{-6}{-2}}\]\[\huge\color{blue}{\frac{3}{1}}~~~~same~~~as~~~3\]

OpenStudy (solomonzelman):

Hold on....

OpenStudy (solomonzelman):

\[\huge-1/\color{blue}{3}\]

OpenStudy (solomonzelman):

\[D\]

OpenStudy (olsolneo123):

thank you!!

OpenStudy (solomonzelman):

Anytime!

OpenStudy (olsolneo123):

you up for more?

OpenStudy (solomonzelman):

On a different thread please...

OpenStudy (solomonzelman):

make a new question (in open questions)

OpenStudy (olsolneo123):

Given the two equations below, determine which statement best describes the graph of the equations. \[y=\left(\begin{matrix}3 \\ 5\end{matrix}\right)x-5\] \[y=-\left(\begin{matrix}5 \\ 3\end{matrix}\right)x\] A.The two lines given by the above equations are perpendicular. B.The two lines given by the above equations intersect (but are not perpendicular). C.The two lines given by the above equations coincide. D.The two lines given by the above equations are parallel.

OpenStudy (solomonzelman):

I think A.

OpenStudy (olsolneo123):

yessssss!

OpenStudy (solomonzelman):

Next time lease post one question per thread.\ I am glad that I helped you though.

OpenStudy (olsolneo123):

ok thank you

OpenStudy (solomonzelman):

Anytme!